8
_____
2+i
I know I need to divide this. The i always throws me off. is it 4 + i?


Answer by stanbon(75887)

8
_____
2+i
---------------
Multiply numerator and denominator by 2-i to get:
= 8(2-i)/(4+1)
= (8/5)(2-i)
= (16/5) - (8/5)i
======================
Cheers,
Stan H.